vivianaXlafuente:

1 gallon = 4 quarts 1 quart = 2 pints 1 pint = 2 cups 1 cup = 8 ounces How many ounces are in 4 gallons? A) 32 ounces B) 64 ounces C) 256 ounces D) 512 ounces

Eiwoh2:

Did you get the answer for this one?

vivianaXlafuente:

no

jhonyy9:

total there are 4 gallons so step by step - us the rule of 3 simple 1 gallon ---- 4 quarts 4 gallon --- x


x = 4*4/1 = 16 so result 4 gallon = 16 quarts 1 quart = 2 pints 16 quarts = x
x= 16*2/1 = 32 so result 16 quarts = 32 pints 1 pint = 2 cups 32 pints = x
x = 32*2/1 = 64 so result 32 pints = 64 cups 1 cup = 8 ounces 64 cups = x
x=64*8/1=512 ounces so from these all result that 4 gallons are 512 ounces hope that is understandably and will help you